OpenStudy (anonymous):

Rewrite the equation 8x+6x^2-7=x^2 in standard form and identify a, b, and c.

OpenStudy (anonymous):

\[8x+6x^2-7=x^2 \rightarrow 8x+6x^2-x^2-7=0\] \[\rightarrow 8x+5x^2-7=0 \rightarrow 5x^2+8x-7=0\] \[5x^2+8x-7=0\] i written in the standard form. Comparing it with \[ax^2+bx+c=0\] we find a= 5, b=8 & c= -7 @xxshaybearxx

OpenStudy (anonymous):

thank you!